### v3

# Opt-Out: Investigating Entity-Level Unlearning for Large Language Models via Optimal Transport

Abstract
Instruction-following large language models (LLMs), such as ChatGPT, have become widely popular among everyday users. However, these models inadvertently disclose private, sensitive information to their users, underscoring the need for machine unlearning techniques to remove selective information from the models. While prior work has focused on forgetting small, random subsets of training data at the instance-level, we argue that real-world scenarios often require the removal of an entire user data, which may require a more careful maneuver. In this study, we explore entity-level unlearning, which aims to erase all knowledge related to a target entity while preserving the remaining model capabilities. To address this, we introduce Opt-Out, an optimal transport-based unlearning method that utilizes the Wasserstein distance from the model’s initial parameters to achieve more effective and fine-grained unlearning. We also present the first Entity-Level Unlearning Dataset (ELUDe) designed to evaluate entity-level unlearning. Our empirical results demonstrate that Opt-Out surpasses existing methods, establishing a new standard for secure and adaptable LLMs that can accommodate user data removal requests without the need for full retraining. [^1] Our code and data are available at https://github.com/brightjade/Opt-Out.
[…]
Opt-Out: Investigating Entity-Level Unlearning for
Large Language Models via Optimal Transport
Minseok Choi       Daniel Rim†, ‡       Dohyun Lee       Jaegul Choo KAIST AI     Hyundai Motor Company {minseok.choi,aiclaudev,jchoo}@kaist.ac.kr drim@hyundai.com
1 Introduction
Machine unlearning (MU) is the task of reversing the learning process that aims to remove the influence of data points from a trained machine learning model. The field has emerged to mitigate the risk of private data leakage upon completion of training Cao and Yang (2015), particularly in compliance with legislations, such as the Right to be Forgotten (RTBF) Rosen (2011) in the European Union’s General Data Protection Regulation (GDPR) Hoofnagle et al. (2019) and the United States’ California Consumer Privacy Act (CCPA) Pardau (2018) requiring the removal of personal information when requested. With research showing that extracting training data becomes easier as large language models (LLMs) scale Carlini et al. (2022), ensuring privacy protections for LLMs has become increasingly crucial.
Figure 1: Motivation for entity-level unlearning. While instance-level unlearning operates at a coarse granularity, entity-level unlearning aims to remove target entity knowledge while carefully preserving knowledge about neighboring entities and the broader world knowledge.
Despite the pressing requirement of the task, eliminating the impact of data samples on billions of model parameters is extremely challenging. The surest approach is exact unlearning, wherein LLMs are completely retrained from scratch using the remaining training set after removing the data points to be forgotten. Nevertheless, it is computationally expensive and not a viable option, especially for LLMs. Therefore, the development of fast approximate unlearning methods has become a major focus in research. Research on MU has primarily been conducted in computer vision tasks Golatkar et al. (2020a, b); Bourtoule et al. (2021); Gandikota et al. (2023); Kurmanji et al. (2023); Fan et al. (2024); however, with the rise of LLMs Brown et al. (2020); Dubey et al. (2024); Abdin et al. (2024), it is gaining prominence in NLP due to privacy problems exhibited by LLMs Zhang et al. (2023).
Recently, several MU approaches in NLP have been proposed Jang et al. (2023); Wang et al. (2023a); Chen and Yang (2023); Lee et al. (2024); Zhang et al. (2024). Notably, Jang et al. (2023) first introduced an unlearning technique that reverses the gradient to prevent LLMs from generating specific sensitive token sequences. However, this often resulted in model collapse, where the model starts to produce low-quality, homogeneous responses, especially as the number of instances to forget increases. To remedy this issue, recent methods have attempted to incorporate additional retention data during training Lee et al. (2024) or to relax the unlearning loss to mitigate collapse Zhang et al. (2024). While these strategies have demonstrated promising results, their evaluations have been limited to small, random sets of instances (i.e., at the instance-levelJang et al. (2023); Maini et al. (2024). Moreover, these methods did not account for a real-world scenario, where a specific person’s data needs to be removed. As illustrated in Figure 1, users may request their personal data be erased under their RTBF. In such cases, it is pivotal to safely and effectively “unlink” the neighboring knowledge while preserving the rest of the information contained in the LLM.
In this work, we investigate entity-level unlearning, which focuses on removing all knowledge associated with a specific entity while retaining the rest of the model’s information. To simulate real-world unlearning scenarios, we introduce the first Entity-Level Unlearning Dataset (ELUDe), consisting of 20 real-world target entities built from their respective Wikipedia pages. Additionally, we create a dataset of 10 neighboring entities for each target, serving as retention data that is closely related to the target entity but should remain unforgotten. To further improve the performance of entity-level unlearning, we propose Opt-Out, a novel fine-grained unlearning method grounded in optimal transport theory. Specifically, Opt-Out employs the Wasserstein distance from the LLM’s initial weights to regularize the unlearning process with the optimal transportation cost between the parameters. This enables fine-grained control over the parameters, maximizing those crucial for unlearning while minimizing those essential for retention. We evaluate our framework on ELUDe, alongside several LLM benchmarks, and demonstrate that Opt-Out outperforms existing unlearning methods in both unlearning and retaining performance, highlighting the effectiveness of our approach. Our work focuses on Wikipedia entities due to their extensive coverage and accessibility, rather than the actual privacy data; however, we hope this work provides a testbed for entity-level unlearning, taking a modest step toward advancing the development of practical unlearning methods.
Figure 2: Overview of our proposed framework, which consists of three key steps: 1) Forget Set Construction, where target entity knowledge is extracted from Wikipedia and GPT-4o is used to create the forget set, covering as much knowledge as it can; 2) Retain Set Construction, following a similar process to build the retain set using knowledge from neighboring entities; and 3) Optimal Transport-based (OT) Unlearning, which computes the Wasserstein distance between two sets of parametric weights and regularizes the model accordingly.
2 Dataset Construction
In this section, we present ELUDe, the first entity-level unlearning dataset focused on the removal of an entire entity. The dataset includes 20 real-world target entities and 144 unique neighboring entities, comprising 15,651 forget samples and 90,954 retain samples. The data collection process is described in detail in the subsequent sections.
2.1 Selecting Target Entities
To reverse the influence of data points on a specific entity, an ideal approach would involve access to the exact subset of data used during pretraining. However, obtaining such data is impractical because the pretraining corpus for most LLMs is often concealed. Even if it were available, isolating the data relevant to a particular entity would be extremely challenging. Therefore, we leverage Wikipedia to extract entity knowledge. Wikipedia serves as a reliable source because its widely recognized information is often memorized by various LLMs, making it suitable for knowledge unlearning. Additionally, previous studies have demonstrated that Wikipedia provides high coverage of information about individuals and maintains reasonable self-consistency Min et al. (2023). We specifically choose 20 target entities from the most popular Wikipedia pages [^2] https://en.wikipedia.org/wiki/Wikipedia:Popular_pages, using page views as a proxy for how frequently these entities are discussed online. For effective unlearning evaluation, we need data that LLMs have heavily memorized. If the model lacks prior knowledge of an entity, assessing unlearning becomes difficult, potentially requiring additional finetuning before unlearning, as observed in Maini et al. (2024). Utilizing popular Wikipedia pages aligns well with LLM memorization and is more cost-effective than exploring large pretraining corpora Mallen et al. (2023).
2.2 Selecting Neighboring Entities
We can employ any kind of textual data for retention, such as TruthfulQA Yao et al. (2023) or Wikitext Li et al. (2024). However, we posit that finetuning with a general retain set may not effectively disentangle entity knowledge from related information. Inspired by hard negatives in representation learning Gillick et al. (2019), we curate the retained data by mining neighboring pages of a given entity. Specifically, for each target entity, we select 10 neighboring entities based on the following criteria: 1) there is a bidirectional relationship, meaning both entities link to each other and are mentioned at least once on their respective pages, 2) the neighboring pages rank within the top 10 in terms of page views over the past three years, and 3) the neighboring pages are all people. For 20 target entities, this process yields 144 unique neighboring entities (due to overlap).
2.3 Generating QA Pairs
After identifying all entities, we transform each corresponding Wikipedia page into a set of QA pairs. While it is technically feasible to input the entire page into an LLM, we find that the QA format works more seamlessly with chat-based models and better simulates real-world interactions. To generate these QA pairs, we process each paragraph through GPT-4o Achiam et al. (2023), prompting it to create as many QA pairs as possible, aiming to cover the full scope of factual content. The specific prompt used for this data generation process is detailed in Appendix F. Since some paragraphs may convey overlapping or identical information, we apply a deduplication step using BERT embeddings from Sentence Transformer Reimers and Gurevych (2019). On average, approximately 647 QA pairs per entity were created.
3 Methodology
3.1 Problem Definition
Knowledge Unlearning
Given a token sequence $\mathbf{x}=\{x\}_{i=1}^{T}$ within the training dataset $\mathcal{D}=\{\mathbf{x}\}_{i=1}^{N}$, the goal of knowledge unlearning is to safely eliminate the influence of a specific subset of data $\mathcal{D}_{f}$ from a trained machine learning model. This process ensures that the model behaves as though the removed data was never included in the training while preserving its performance on the remaining dataset. Conventionally, the data to be forgotten $\mathcal{D}_{f}$ is referred to as the forget set, and the data to be retained is called the retain set. For simplicity, we focus on the standard scenario where $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$ are mutually exclusive (i.e., $\mathcal{D}_{f}\cap\mathcal{D}_{r}=\varnothing$).
Entity-Level Unlearning
Entity-level unlearning aims to remove all knowledge related to a specific target entity from a trained machine learning model. However, because the knowledge about an entity within a model is inherently abstract and cannot be explicitly predefined, we adopt the framework proposed by Ma et al. (2024). This reformulates the task as erasing all knowledge associated with a target entity by identifying and deleting a specific subset of entity-associated knowledge, which we term the “proxy forget set”. The proxy forget set encapsulates identifiable aspects of the target entity’s knowledge. Since it is equally impractical to obtain the exact retain set used during pretraining, we define the “proxy retain set” as a specific subset of knowledge encompassing everything except the target entity.
Formally, for a target entity $t$ and its proxy forget set $\mathcal{D}_{f}^{t}$, the objective of entity-level unlearning is to train a model $\phi_{\theta}$ such that the updated model $\phi_{\theta^{\prime}}=S(\phi_{\theta};\mathcal{D}_{f}^{t})$ operates as if it were trained exclusively on the proxy retain set $\mathcal{D}_{r}^{t}$. The unlearning function $S$ ensures that $\mathcal{D}_{f}^{t}$ is effectively forgotten while retaining the model’s performance on $\mathcal{D}_{r}^{t}$. It is important to note that the exact forget and retain sets are unattainable, except in highly controlled scenarios such as fictitious unlearning Maini et al. (2024). Consequently, achieving high knowledge coverage in the proxy sets directly improves alignment with the ideal performance attainable using exact sets. To address this, we leverage Wikipedia as our knowledge source due to its extensive coverage of entity-specific information, thereby narrowing the gap between proxy and exact sets.
3.2 Knowledge Unlearning
The primary goal of language modeling is to minimize the negative log-likelihood of token sequences, training the model to accurately predict the next token in a sequence. To remove specific knowledge from language models, a straightforward approach is to apply gradient ascent on the next-token prediction loss over the forget set, which can be understood as equivalent to gradient descent on the negative prediction loss:
$\mathcal{L}_{\text{GA}}=-\mathbb{E}_{\mathcal{D}_{f}^{t}}[-\log(\phi_{\theta}(y|x))].$ (1)
By inverting the language modeling objective, many existing unlearning methods have successfully removed parametric knowledge of the forget set from language models. However, numerous studies have highlighted the catastrophic effects of gradient ascent Yao et al. (2023); Lee et al. (2024). To address these issues, Zhang et al. (2024) introduced negative preference optimization (NPO), a technique that simplifies to gradient ascent in the high-temperature limit but is inherently more stable and lower-bounded, significantly slowing the model collapse compared to gradient ascent. NPO draws on preference optimization Rafailov et al. (2023) and aligns the language model with negative examples exclusively:
$\mathcal{L}_{\text{NPO}}=-\mathbb{E}_{\mathcal{D}_{f}^{t}}\left[\log\sigma\left(-\eta\log\frac{\phi_{\theta}(y|x)}{\phi_{\text{ref}}(y|x)}\right)\right],$ (2)
where $\sigma$ represents the sigmoid function, $\eta>0$ is the inverse temperature, and $\phi_{\text{ref}}$ is a reference model. In entity-level unlearning, we observe that the NPO loss also produces much more stable and reliable results in practice. However, finetuning solely on the forget set eventually leads to model degradation and collapse. As with prior unlearning methods, we also train the model on the retain set to explicitly preserve the remaining knowledge. This is achieved through standard language modeling on the retain set, which serves as the positive counterpart to Equation 1: $\mathcal{L}_{\text{RT}}=-\mathbb{E}_{\mathcal{D}_{r}^{t}}[\log(\phi_{\theta}(y|x))]$.
3.3 Optimal Transport-Based Unlearning
To further enhance the performance of entity-level unlearning, we propose Opt-Out, a fine-grained unlearning approach grounded in optimal transport theory. Building on this theory, we develop the Wasserstein regularization, which calculates the Wasserstein distance between two sets of parametric weights and regularizes the model based on this distance. The Wasserstein distance, also known as Earth Mover’s Distance, addresses the optimal transport problem by measuring the minimum effort required to move one distribution of mass to another. We hypothesize that computing this distance helps us estimate the optimal transportation cost between parameters, facilitating more effective unlearning. By applying this framework, we allow more significant shifts in parameters that are crucial for unlearning, while reducing changes in parameters important for retention. In mathematical terms, given a source distribution $\mu$ and a target distribution $\nu$, sampled from probability space $\mathbb{X},\mathbb{Y}\in\Omega$ respectively, the optimal transport attempts to compute the minimal transportation cost between the two distributions. Formally, Kantorovich (2006) formulates the problem with a probabilistic coupling $\pi\in\mathcal{P}(\mathbb{X}\times\mathbb{Y})$:
[…]
$\begin{split}\bm{\gamma}^{*}&=\operatorname*{arg\,min}_{\gamma\in\mathbb{R}_{+}^{m\times n}}\sum_{i=1}^{m}\sum_{j=1}^{n}\gamma_{ij}C_{ij}\\ &\text{s.t. }\gamma\bm{1}=\alpha,\gamma^{\top}\bm{1}=\beta,\gamma\geq 0,\end{split}$ (4)
where $\bm{\gamma}^{*}$ is the optimal transport plan or transport matrix, $C\in\mathbb{R}_{+}^{m\times n}$ is the cost matrix defining the cost to move mass from bin $\alpha_{i}$ to bin $\beta_{j}$, and $\alpha$ and $\beta$ are histograms on the simplex that represent the weights of each sample in the source and target distributions. Building on the optimal transport equation, given the initial weights of the language model as $\theta_{0}$, the Wasserstein distance between $\theta_{0}$ and the training parameters $\theta$ with finite $p$-moments is then computed as
$\begin{split}W_{p}(\theta,\theta_{0})&=(\min_{\gamma\in\mathbb{R}_{+}^{m\times n}}\sum_{i,j}\gamma_{ij}||\theta_{i}-\theta_{0,j}||_{p})^{\frac{1}{p}}\\ &\text{s.t. }\gamma\bm{1}=\alpha,\gamma^{\top}\bm{1}=\beta,\gamma\geq 0.\end{split}$ (5)
However, it is intractable to compute the exact $\bm{\gamma}^{*}$, because the time complexity of the exact solver is $O(n^{3}\log n)$ and the memory complexity is always $O(n^{2})$ due to the cost matrix. Especially for LLMs, the number of parameters exceeds billions, if not trillions. For efficiency in both time and memory, we approximate the Wasserstein distance by computing the Sliced Wasserstein Distance (SWD) Bonneel et al. (2015). Instead of computing the entire cost matrix, SWD reduces the dimensionality of the problem by projecting the distributions onto random slices and then computing the Wasserstein distance in a lower-dimensional space. Concretely, the Monte Carlo approximation of the $p$-sliced Wasserstein distance is given by
$SW_{p}(\theta,\theta_{0})=\mathop{{}\mathbb{E}}_{u\sim\mathcal{U}(\mathbb{S}^{d-1})}(W_{p}(u_{\#\theta},u_{\#\theta_{0}}))^{\frac{1}{p}},$ (6)
where $\mathcal{U}(\mathbb{S}^{d-1})$ denotes the uniform distribution on the unit sphere in $\mathbb{R}^{d}$, and $u_{\#\theta}$ and $u_{\#\theta_{0}}$ stand for the pushforwards of the projections of $\theta$ and $\theta_{0}$ along the direction of $u\in\mathbb{S}^{d-1}$, respectively. Putting everything together, the overall training objective for fine-grained entity-level unlearning is minimizing the following loss:
$\mathcal{L}=\mathcal{L}_{\text{NPO}}+\mathcal{L}_{\text{RT}}+\lambda\cdot SW_{p}(\theta,\theta_{0}),$ (7)
where $\lambda$ is a hyperparameter for scaling the regularization term.
FQ RQ ARC-C CSQA Hella. Lamba. MMLU OBQA PIQA Wino. Avg.
Llama-3.1-8B-Instruct
Original 45.5 51.2 51.8 77.1 59.2 73.2 68.1 33.8 80.2 74.1 64.7
Guardrail 64.8 51.3 49.8 75.5 58.9 72.3 67.2 33.2 80.1 74.2 63.9
GA* 70.9 0.0 23.6 21.9 32.2 11.6 33.9 28.0 58.3 60.9 33.8
DPO* 76.3 0.0 22.9 30.5 52.8 36.4 37.9 28.3 58.4 69.5 42.1
NPO* 89.7 0.0 24.7 23.0 37.6 20.8 36.3 30.6 60.1 67.2 37.5
IDK* 84.3 3.5 38.4 67.6 53.3 48.0 61.8 30.0 77.1 71.3 56.0
GA+RT 77.1 45.7 47.4 71.0 57.7 71.1 60.7 32.9 79.2 72.2 61.5
[…]
Phi-3.5-Mini-Instruct
Original 44.9 34.9 59.5 75.3 58.8 65.1 68.7 37.6 80.0 74.6 65.0
Guardrail 46.6 26.4 52.2 72.9 59.2 64.1 67.7 38.3 78.1 74.5 63.4
GA* 63.6 0.0 24.0 19.8 52.1 47.0 23.6 32.6 53.8 66.0 39.8
DPO* 78.3 0.0 38.9 36.1 56.7 54.3 54.0 36.4 63.9 72.7 51.6
NPO* 80.7 0.0 28.6 19.6 56.8 57.0 26.6 35.8 59.1 69.4 44.1
IDK* 80.4 4.3 53.7 70.9 54.8 47.8 65.7 36.6 79.4 76.7 60.7
GA+RT 67.7 47.3 56.9 69.4 56.7 56.8 67.2 35.8 79.9 73.1 62.0
[…]
Table 1: Performance (%) of various methods after unlearning on Llama-3.1-8B-Instruct and Phi-3.5-Mini-Instruct. FQ (Forget Quality) reflects the harmonic mean of ground-truth token probabilities, ROUGE-L recall scores, and truth ratio over the forget set, while RQ (Retain Quality) is computed across the retain and world sets. Methods are also assessed on eight LLM benchmarks to evaluate the retention of overall model capabilities. (*) indicates collapsed models. The best results are in bold, while the second best are underlined.
4 Experiments
4.1 Datasets
We utilize the forget and retain sets from ELUDe to evaluate entity-level unlearning. The retain set is divided into training, validation, and test splits in an 8:1:1 ratio. Since the training portion of the retained data is significantly larger than the forget set, we apply random sampling during training. Moreover, we incorporate the Alpaca-GPT4 instruction dataset Peng et al. (2023) as an auxiliary retain set (i.e., world set) to align the model with general instructional tasks. Specifically, we use 50k instructional examples for training, 1k for validation, and 1k for testing. To assess model utility, we also validate our framework on eight language understanding benchmarks including ARC-Challenge Clark et al. (2018), CommonsenseQA Talmor et al. (2019), HellaSwag Zellers et al. (2019), Lambada Paperno et al. (2016), MMLU Hendrycks et al. (2021), OpenbookQA Mihaylov et al. (2018), PIQA Bisk et al. (2020), and Winogrande Sakaguchi et al. (2021).
4.2 Evaluation Metrics
Following closely with Maini et al. (2024), we measure the unlearning performance using a stack of the following metrics:
Probability
We compute the conditional probability $P(a|q)$ for the forget and retain sets, normalizing for answer length by raising it to the power $1/|a|$. For the world set, each question $q$ is treated as multiple-choice with choices $\{a_{1},...,a_{n}\}$, where $a_{1}$ is the correct answer. The probability is then $P(a_{1}|q)/\sum_{i=1}^{n}P(a_{i}|q)$.
ROUGE
We use ROUGE-L recall Lin (2004) to compare model answers (greedy sampling) with ground truth, serving as a proxy for QA accuracy by accounting for variations in phrasing.
Truth Ratio
We compute a ratio comparing the likelihood of the correct answer to incorrect ones. Since finetuning may inflate the probability of the exact ground truth phrasing, we use a paraphrased version of the correct answer and average probabilities over multiple similarly formatted wrong answers. This ratio helps assess whether the unlearning algorithm removed the target information, even if the model no longer provides exact matches but still favors correct responses. Let $\tilde{a}$ denote the paraphrased answer and $\mathcal{A}_{\text{pert}}$ denote a set of five perturbations generated by GPT-4o. The truth ratio $R_{\text{truth}}$ is given by:
$R_{\text{truth}}=\frac{\frac{1}{|\mathcal{A}_{\text{pert}}|}\sum_{\hat{a}\in\mathcal{A}_{\text{pert}}}P(\hat{a}|q)^{q/|\hat{a}|}}{P(\tilde{a}|q)^{q/|\tilde{a}|}}$ (8)
For Forget Quality (FQ), we compute the harmonic mean of the three values on the forget set [^3] Unlike in Maini et al. (2024), we do not use the $p$-value from the Kolmogorov-Smirnov test as FQ because it is impossible to compare against a perfectly unlearned model in our setup, and even more so in real-world applications. For comparison with exact unlearning, refer to Appendix E for our results on the TOFU dataset., while for Retain Quality (RQ), we take the harmonic mean of the six values across both the retain and world sets to prevent low scores from getting averaged out. Some values are inverted so that higher values indicate better performance (e.g., $\text{max}(0,1-R_{\text{truth}})$ is used in RQ).
Figure 3: Forget Quality performance (%) of different +RT methods following unlearning on Llama-3.1-8B-Instruct, evaluated against nine types of adversarial prompt attacks. Each attack is described in detail in Appendix G.
4.3 Baselines
We compare our framework with the following unlearning methods:
  • Guardrail Thaker et al. (2024): A simple prompting baseline that instructs the LLM to refuse to answer about the specified entity
  • GA Jang et al. (2023): Applies gradient ascent on the forget set
  • DPO Rafailov et al. (2023): Employs direct preference optimization where “I don’t know” responses are preferred on the forget set
  • NPO Zhang et al. (2024): Utilizes negative preference optimization on the forget set
  • IDK Maini et al. (2024): Finetunes the model to provide “I don’t know” responses for the forget set
  • +RT: Additionally finetunes the model on the retain set for explicit model retention
4.4 Unlearning Results
We present a comparison of unlearning results across various methods in Table 1. Our experiments follow the single-target unlearning setting, where one target is forgotten at a time, with the results averaged over five unlearning targets. First, we observe that Guardrail, which utilizes the system prompt “If the question asks about {entity}, say you do not know the answer; otherwise, answer as best as you can,” effectively retains information but struggles to adequately forget the target entity. For the Phi-3.5 model, Guardrail negatively impacts RQ performance, indicating that in-context unlearning is not suitable for smaller models. Unlearning baselines such as GA, DPO, NPO, and IDK show improvements in FQ; however, these methods tend to collapse, with RQ dropping to near zero and overall benchmark performance significantly degrading. With additional finetuning on the retain set (+RT), retention performance improves across the board, while FQ remains strong. Notably, Opt-Out outperforms all methods across both Llama-3.1 and Phi-3.5 models and maintains competitive RQ and overall LLM benchmark performance, demonstrating the effectiveness of our proposed approach.
4.5 Performance Against LLM Attacks
Membership Inference Attacks
We assess performance against Membership Inference Attacks (MIAs) to ensure that, after unlearning, an attacker cannot distinguish between unlearned examples and those never seen by the model, thus protecting user privacy. Following Chen and Yang (2023), we train a binary classifier (the “attacker”) on the unlearned model’s losses for forget and test samples. Since we perform entity-level unlearning on the entire forget set, we use a paraphrased set for the test samples. Ideally, 50% accuracy indicates the attacker cannot differentiate between the two, validating the unlearning method. As shown in Table 2, most unlearned models, including Opt-Out, successfully defend against MIAs.
Llama-3.1 Phi-3.5
Method mean std mean std
Oracle 50.0 - 50.0 -
GA 53.8 2.9 54.7 1.4
DPO 56.2 3.8 53.8 2.2
NPO 53.8 1.8 52.8 1.6
IDK 59.3 2.3 58.7 1.8
GA+RT 50.9 3.1 49.7 1.0
DPO+RT 50.3 3.1 49.7 3.1
NPO+RT 49.6 2.5 50.9 2.5
IDK+RT 69.1 1.5 67.3 1.9
Opt-Out (ours) 48.6 1.0 49.1 1.1
Table 2: MIA accuracy (%) of a trained binary classifier (“the attacker”) predicting whether an input data belongs to the training set. 50% indicates the best performance.
Adversarial Prompt Attacks
Given the use of instruction-tuned models, safeguarding against malicious prompt attacks is vital. To rigorously evaluate the efficacy of unlearning in mitigating adversarial attacks, we follow Jin et al. (2024) and assess unlearned models against nine different types of adversarial threats. Detailed descriptions of the attack examples are provided in Appendix G. As illustrated in Figure 3, our proposed approach, Opt-Out, consistently achieves high-quality forgetting across various adversarial attacks, demonstrating strong robustness against malicious prompts.
4.6 Effect of Wasserstein Regularization
We verify the effectiveness of the proposed Wasserstein regularization by comparing it to other commonly used distance metrics. As shown in Table 3, the Manhattan distance preserves the most information, but this is largely attributed to the fact that the model underwent minimal unlearning due to excessively strong regularization. In contrast, the Euclidean and Cosine distances show reasonable unlearning performance, though they slightly underperform compared to using no regularization at all (as evidenced by NPO+RT in Table 1). In comparison, our proposed Wasserstein distance delivers the best overall results, highlighting the efficacy of optimal transport-based unlearning.
Distance Metric FQ RQ Util. Wasserstein (ours) 87.8 46.6 63.3 Manhattan 47.0 50.9 64.6 Euclidean 81.5 46.2 63.0 Chebyshev 86.3 45.4 62.2 Cosine 81.6 45.8 62.8
Table 3: Comparison of distance metrics in regularization with Llama-3.1-8B-Instruct. Util. is the average of results across the eight LLM benchmarks.
4.7 Effect of Neighboring Entity Data
To validate the effectiveness of our neighboring entity data augmentation, we measure the unlearning performance of a model trained without the neighboring entity set, using only the world set (i.e., Alpaca-GPT4). As illustrated in Figure 4, the model trained solely on the world set shows comparable performance in terms of Forget Quality and overall model utility but exhibits significantly worse performance on Retain Quality. We attribute this to the model’s difficulty in distinguishing between forget and retain examples when trained exclusively on world data. In contrast, the model supplemented with our neighboring entity data consistently outperforms the other settings across all metrics, highlighting the importance of incorporating closely related data, which likely acts as “hard positives,” aiding the model in better differentiating forget and retain examples.
Figure 4: Performance comparison between using only the world set and supplementing it with our neighboring entity set as retain data during training. Scores are averaged across GA, DPO, NPO, IDK, and Opt-Out methods using Llama-3.1-8B-Instruct.
5 Related Work
5.1 Machine Unlearning
With the emergence of machine unlearning to mitigate privacy concerns Cao and Yang (2015); Golatkar et al. (2020a); Kurmanji et al. (2023), the focus of unlearning techniques in computer vision has predominantly centered on image classification models where they aim to forget a whole class, thereby attaining random performance for particular image classes. Recently, there have been attempts to perform unlearning in image generation Fan et al. (2024) or erase specific concepts from diffusion model weights, utilizing negative guidance as a teacher to drive the unlearning process Gandikota et al. (2023). Concept erasure aims to identify and remove specific concepts that may be encoded  Ravfogel et al. (2022a, b); Belrose et al. (2023), applying various transformations to the neural representations. These methods generally approach the problem from a theoretical setting and look to identify and erase a high-level concept that may cause biases, such as gender or racial biases.
5.2 Knowledge Unlearning
Likewise, the primary emphasis of unlearning in NLP has been directed towards tasks such as text classification and generation Wang et al. (2023a); Chen and Yang (2023); Yao et al. (2023). Introducing a new paradigm, Jang et al. (2023) proposed unlearning specific token sequences by negating the gradient descent. Nevertheless, this often led to model collapse, especially as the number of samples to forget increased. To address this issue, Lee et al. (2024) presented a more robust method to mitigate performance degradation by incorporating retention mechanisms. Others shared similar concerns about catastrophic failure in machine unlearning and suggested solutions based on preference optimization Zhang et al. (2024). These methods, however, primarily target unlearning specific instances in language models. In this work, we focus on removing targeted entity-level information that may have been learned during pretraining, leveraging an optimal transport-based technique for more effective and fine-grained unlearning. A concurrent work Ma et al. (2024) also explores entity-level unlearning but is limited to the task of fictitious unlearning Maini et al. (2024).
5.3 Unlearning Datasets
With the latest development of machine unlearning for LLMs, the need for dedicated unlearning datasets and benchmarks has become increasingly important. Li et al. (2024) introduced the Weapons of Mass Destruction Proxy (WMDP) benchmark, which includes 3,668 multiple-choice questions designed to measure hazardous knowledge in biosecurity, cybersecurity, and chemical security. Maini et al. (2024) presented the Task of Fictitious Unlearning (TOFU), featuring 20 QA pairs for each of 200 fictitious authors. Jin et al. (2024) released the Real-World Knowledge Unlearning (RWKU) benchmark, focusing on 200 real-world celebrities and comprising 2,879 QA pairs. In parallel, our work introduces a new dataset ELUDe, which includes 20 real-world popular entities. Unlike previous efforts, we provide a substantial volume of data for each entity, totaling 15,651 and 90,954 QA pairs for forget and retain samples, respectively. This enables the complete removal of all knowledge associated with a specific entity, providing a valuable resource for researchers and practitioners tackling real-world user unlearning requests.
6 Conclusion
In this work, we explore entity-level unlearning, a pivotal and timely technique for removing a specific person’s data from LLMs. To simulate real-world user unlearning requests, we introduce ELUDe, a QA dataset designed to train LLMs to selectively forget a specific entity. Furthermore, we propose Opt-Out, an optimal transport-based unlearning method that applies Wasserstein regularization to the model parameters. Our approach outperforms existing unlearning techniques, likely due to its more fine-grained control in knowledge unlearning. These findings are particularly relevant for LLMs deployed in real-world scenarios, enabling them to handle user requests to remove personal data without the need for full retraining.
Limitations
While our framework shows promising performance in unlearning entity-level knowledge, several areas warrant further refinement. First, our work focuses on unlearning Wikipedia entities, which may differ slightly from erasing data related to actual users. Nevertheless, creating meaningful forget and retain sets for an arbitrary person (e.g., Alice) is challenging, as it is difficult to capture how much the LLM knows about her. Therefore, we have leveraged Wikipedia, where the pages themselves serve as a useful proxy for comprehensive data coverage of a particular entity, enabling effective evaluation of full entity-level erasure. Future work could extend our approach to real-world privacy data, incorporating advanced anonymization techniques to better align with practical use cases. Second, our method remains susceptible to generating gibberish post-unlearning. Although it effectively removes parametric knowledge, ensuring the LLM functions correctly for a seamless end-user experience in real-world deployment remains an issue. Combining with the IDK method or remapping outputs to automated responses after unlearning could be considered a simple fix. Lastly, due to computational constraints, we were unable to test models at the scale of 70B parameters or larger. Exploring unlearning techniques with much larger models would better align with the behavior of proprietary models.
Acknowledgements
We are deeply grateful to Hojoon Lee and Dongyoon Hwang for generously sharing their expertise during our discussions. We also appreciate the thoughtful feedback provided by the anonymous reviewers. This work was supported by the Institute for Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program (KAIST), RS-2025-02304967, AI Star Fellowship (KAIST)), and Samsung Electronics Co., Ltd.
Appendix A Implementation Details
Our framework is built on PyTorch Paszke et al. (2019), Hugging Face Transformers Wolf et al. (2020), and Accelerate Gugger et al. (2022). We use Llama-3.1-8B-Instruct Dubey et al. (2024) and Phi-3.5-Mini-Instruct Abdin et al. (2024) and optimize their weights with AdamW Loshchilov and Hutter (2019), tuning hyperparameters to maximize Forget and Retain Quality. We set the batch size to 32, the learning rate to 1e-5, the weight decay to 0.01, the inverse temperature $\eta$ to 0.1, and the regularization strength $\lambda$ to 0.1. We train for 3 epochs and use early stopping if the model performance decreases from the last epoch. All experiments are conducted with four NVIDIA H100 GPUs.
Appendix B Human Evaluation
To verify the reliability of our machine-generated dataset, we perform human evaluation based on the following criteria (0-1 scale):
  1. 1.
    Relevance: Does the question discuss the entity (1 if it does, 0 if not)?
  2. 2.
    Diversity: Is there a similar question in the dataset (0 if there is, 1 if not)?
  3. 3.
    Factuality: Does the answer match with the passage (1 if correct, 0 if it’s hallucinated)?
Following Wang et al. (2023b), we asked authors of this paper to judge training instances for a particular entity on both forget and retain sets. Due to the substantial size of the retain set, we match its number with the corresponding forget set. The evaluators coordinated the standards before starting annotation and then each of them rated all the instances independently. Table 4 shows the average scores for each criterion. We notice that GPT-4o is highly capable of generating relevant, diverse, and factually accurate QA pairs based on the given passage. However, since we feed GPT-4o one passage at a time, some facts tend to overlap with those from previous passages. Additionally, our prompting approach, which encourages GPT-4o to include as many factual details as possible, often results in QA pairs that feature information either not directly related to the main entity (e.g., “What is Cristiano Ronaldo’s mother’s occupation?”) or trivial (e.g., “Which national team does Ronaldo play for?”). The generated QA pairs were predominantly accurate, though any minor factual discrepancies likely stemmed from Wikipedia’s frequent updates. The evaluation results for the retain set were relatively high due to the involvement of multiple entities, which made fact overlap less likely. Moreover, examples were considered acceptable as long as they did not discuss the primary target entity.
Forget Set Retain Set Relevance 89.0 96.0 Diversity 90.0 99.5 Factuality 99.0 98.5
Table 4: Human evaluation results (%).
Appendix C Computational Cost Analysis
To portray the feasibility of Opt-Out, we conduct a computational cost analysis and present the results in Table 5. Opt-Out introduces a small additional overhead during training while maintaining an identical inference cost to existing methods, ensuring its practicality for large-scale applications. Specifically, the additional overhead stems from computing the Sliced Wasserstein Distance (SWD) between two sets of parametric weights. Regarding memory usage during training, Opt-Out is comparable to DPO and NPO, which also require holding an extra reference (frozen) model. The additional memory cost in Opt-Out arises from a lightweight random projection layer used in SWD computation.
Given that $L$ is the input sequence length, the training time and memory complexities of Transformer is $O(L^{2})$. +RT incurs additional overhead due to an extra forward pass, while DPO/NPO require a third forward pass through a reference model (DPO demands extra two passes for the preference loss). Opt-Out adds a small training overhead of $O(L\log L)$ for SWD computation (due to sorting) and $O(Ld)$ for projected representations, where $d$ is the dimension of the random projection layer and is kept small. We believe these additional costs are manageable even at scale, as the $L\log L$ term grows significantly slower compared to the $L^{2}$ complexity of Transformer training itself.
Training Inference
Method Time Memory Time Memory
GA+RT $\mathcal{O}(2L^{2})$ $\mathcal{O}(L^{2})$ $\mathcal{O}(L)$ $\mathcal{O}(L^{2})$
DPO+RT $\mathcal{O}(5L^{2})$ $\mathcal{O}(2L^{2})$ $\mathcal{O}(L)$ $\mathcal{O}(L^{2})$
NPO+RT $\mathcal{O}(3L^{2})$ $\mathcal{O}(2L^{2})$ $\mathcal{O}(L)$ $\mathcal{O}(L^{2})$
IDK+RT $\mathcal{O}(2L^{2})$ $\mathcal{O}(L^{2})$ $\mathcal{O}(L)$ $\mathcal{O}(L^{2})$
Opt-Out $\mathcal{O}(3L^{2}+L\log L)$ $\mathcal{O}(2L^{2}+Ld)$ $\mathcal{O}(L)$ $\mathcal{O}(L^{2})$
Table 5: Computational complexity of various methods (constants included to better depict the comparison). Opt-Out incurs a slight additional overhead during training while maintaining an identical inference cost to existing methods, ensuring its practicality.
Forget Set Retain Set World Set
Prob.($\downarrow$) ROUGE($\downarrow$) TR($\uparrow$) Prob.($\uparrow$) ROUGE($\uparrow$) TR($\uparrow$) Prob.($\uparrow$) ROUGE($\uparrow$) TR($\uparrow$)
Llama-3.1-8B-Instruct
Original 40.7 63.7 46.4 38.6 61.4 50.7 53.1 47.7 64.8
Guardrail 26.5 13.5 47.4 38.7 62.0 50.6 53.5 47.1 64.9
GA* 0.0 0.0 44.8 0.0 0.0 21.3 0.0 0.2 37.0
DPO* 0.0 1.1 52.1 0.0 1.1 20.6 0.0 1.2 60.4
NPO* 0.0 0.0 74.3 0.0 0.0 10.2 0.0 0.2 31.9
IDK* 10.9 1.0 70.0 13.3 1.0 31.1 44.6 1.7 56.1
[…]
Phi-3.5-Mini-Instruct
Original 11.1 64.6 36.6 11.3 63.3 61.1 58.1 50.3 71.5
Guardrail 8.2 57.2 33.2 6.9 55.8 64.2 60.4 50.4 74.2
GA* 0.0 0.1 36.9 0.0 0.2 27.4 0.0 2.0 51.0
DPO* 0.0 0.9 54.9 0.0 0.8 19.6 0.0 1.7 56.6
NPO* 0.0 0.1 58.2 0.0 0.2 19.9 0.0 1.5 53.8
IDK* 22.1 1.1 69.7 23.2 1.0 31.5 41.2 3.5 56.6
[…]
Table 6: Detailed unlearning results on Llama-3.1-8B-Instruct and Phi-3.5-Mini-Instruct.
Appendix D Full Evaluation Results
[…]
Appendix E Exact Unlearning Results
We conduct additional experiments on the TOFU dataset Maini et al. (2024) to further demonstrate the effectiveness of our method and report the results in Table 7. The TOFU dataset is entirely fictitious and was not included during the model’s pretraining. This ensures that the model starts as “perfectly unlearned”, enabling a direct comparison between our approach and exact unlearning. Forget Quality (FQ) in this context is calculated using $p$-values derived from the Kolmogorov-Smirnov test (KS-Test), which quantifies the divergence between the distributions of Truth Ratios in unlearned models and the exact unlearning model. Retain Quality (RQ) is also slightly adapted here, reflecting the harmonic mean of nine values (Probability, ROUGE, and Truth Ratio measured across Retain Set, Real Authors, and World Facts). For clarity, an asterisk (*) differentiates terms redefined in this context from those in our main paper. The results demonstrate that Opt-Out consistently outperforms other approaches, highlighting its robustness.
Appendix F Prompts
We display the prompt templates used to generate QA pairs for ELUDe in Figure 5, as well as paraphrased and perturbed QA pairs for the truth ratio evaluation in Figures 6 and 7.
Llama-3.1 Phi-3.5
Method FQ* RQ* FQ* RQ*
GA+RT 0.0 16.6 46.6 45.2
DPO+RT 0.0 47.7 1.6 44.7
NPO+RT 11.2 46.1 71.3 45.9
IDK+RT 0.0 41.0 22.1 39.3
Opt-Out (ours) 86.6 40.6 86.6 45.9
Table 7: Performance (%) of various methods after unlearning fictitious data in TOFU (5% forget set setting). FQ* is calculated using $p$-values derived from the KS-Test, and RQ* reflects the harmonic mean of nine values (three metrics measured across three subsets in TOFU).
Appendix G Dataset Examples
We exhibit dataset examples for one of the target entities Cristiano Ronaldo in Figure 8. For the IDK method, we randomly sample from 100 “I don’t know” (IDK) responses in Maini et al. (2024) and replace it with the original response, as shown in Figure 9. To generate adversarial attack prompts, we slightly modify the prompt template used in Jin et al. (2024) to synthesize nine types of adversarial prompt attacks given the original QA pair. We employ GPT-4o to generate 100 examples for each type, making it a total of 900 attack samples for each entity. Examples for each attack type are illustrated in Figure 10.
Figure 5: Prompt template for generating QA pairs for target and neighboring entities.
Figure 6: Prompt template for generating paraphrased QA pairs for evaluation.
Figure 7: Prompt template for generating perturbed QA pairs for evaluation.
Figure 8: Dataset example for the target entity Cristiano Ronaldo. Only the first Wikipedia passage and the first few QA pairs are shown for brevity.
Figure 9: Dataset example for the target entity Cristiano Ronaldo with “I don’t know” responses.
Figure 10: Adversarial prompt attack examples for the target entity Cristiano Ronaldo.